(a) Find the net electrostatic force at sphere B. (b) Find the net electric field at (4,-3)


1
Expert's answer
2022-04-06T14:02:39-0400


Electrostatic force

"F_B=\frac{kq_Aq_B}{r^2}"

"R=\sqrt{4^2+(-3)^2}=5m"

"F_B=\frac{kq_Aq_B}{5^2}=\frac{kq_Aq_B}{25}"

"E_A=\frac{F_B}{q_B}=\frac{kq_A}{5^2}=\frac{kq_A}{25}"